Using the Clausius-Clapeyron equation,calculate the heat of vaporization for the vapour pressure of water at 85 degrees celcius is 0.5atm and at 95 degrees celcius is at 0.9atm.


1
Expert's answer
2021-05-12T06:25:02-0400



"ln(\frac{P_1}{P_2})=(\frac{H_{vap}}{R})(\frac{1}{T_2}-\frac{1}{T_1})"


"ln(\frac{0.5}{0.9})=(\frac{H_{vap}}{R})(\frac{1}{368}-\frac{1}{358})"


"(\frac{H_{vap}}{R})= 3363.06"


"(H_{vap})= 3363.06\u00d78.3145=27962.16 J mol"